# Steering Large Language Models for Machine Translation Personalization

### v1

@@ Appendix B Experiments reproducibility
B.1 Base prompt
We use the same prompt template across all methods: ZS (which corresponds to the original model translation), ZS-Expht, ZS-Exppt (detailed in Section B.5), MS, ActAdd, ReFT, and SAE-based contrastive setups. This prompt, shown in Listing LABEL:lst:explain-final, instructs the model to translate the source sentence while explicitly preventing it from adding any explanations about the translation process. Since all test models are Instruction Tuned, we utilize their native chat templates to preprocess the input accordingly. For multi-shot examples, the user and assistant turns are repeated for each example, always using the same prompt structure.
B.2 Classifier training
@@
Training data only includes generations from models and the translator without any source text. It is also perfectly balanced, as each paragraph provides one instance for all three labels: H1, H2, and MT. The total size of the training set varies depending on the number of paragraphs in the chosen novel. On average, we obtain approximately 830 instances, resulting in a total of around 2,490 labeled examples for training (see 5). Validation and test sets are strictly held out and never seen during training. Additionally, they do not include the small 20-example subsets used in the MS, ZS-Expht, ZS-Exppt, SAE Cont.ht, and SAE Cont.pt setups.
B.3 ReFT training
ReFT training was conducted using the PyReFT toolkit from the original authors141414stanfordnlp/pyreft. We applied the intervention at the same hook point used by other steering methods - specifically, the layer output corresponding to the residual stream at the selected layer. The training configuration includes a low rank dimension of 4, lora alpha set to 32, and a lora dropout of 0.05. ReFT was trained on the same 20 prompts used in the MS setup, for a total of 100 epochs.
B.4 SAE Cont.ht and SAE Cont.pt
@@ B.3 ReFT training
B.5 ZS-Expht and ZS-Exppt
For both the ZS-Expht and ZS-Exppt setups, we used GPT-4o (June 2025) to generate explanations detailing the stylistic differences between a base translation and a target human translation. The prompt template used for this task is shown in Listing LABEL:lst:explain-template, using the same 20 examples as in the MS, SAE Cont.ht, and SAE Cont.pt setups.
All outputs were manually inspected to ensure no verbatim excerpts from the provided examples were present, avoiding any risk of data leakage. Example outputs for different novels are shown in Listing LABEL:lst:explain-example.
Finally, these generated guidelines are used to prompt the evaluated models, following the template shown in Listing LABEL:lst:explain-final.
@@ B.5 ZS-Expht and ZS-Exppt
Base translation: <MT / H- text here>
@@
Appendix C All models results
C.1 Classifiers
@@
Figure 5: Probing classifier performance on the human translation detection task accross Gemma 2 9B layers. For our experiments we select layer 21 as the optimal intervention point for our steering approaches.
@@
C.2 Probing experiments
We report in Figure 5 the probing classifier performance for Gemma 2 9B. We select the 21th layer as the most informative one to make our intervention. Results for the smaller model Gemma 2 2B are shown in the main body of the paper in Figure 2.
C.3 Prompting and steering results
@@
Appendix D Examples from dataset and different approaches
We present in Tables 7 and 8 a selection of examples from two different languages, showcasing outputs from each of the tested setups. For each example, we also report the corresponding classification label predicted by the [Uncaptioned image] classifier and the associated [Uncaptioned image] Comet score. Additionly table 9 shows some examples of models generating output aligned with the Human translator according to the [Uncaptioned image] classifier but with a low [Uncaptioned image] Comet score corresponding to an almost unreadable output due to extreme $\alpha$ values.
@@ Appendix D Examples from dataset and different approaches
ZS-Expht And as he turned his gaze back upon the scene once more, there was displayed an archer’s bow strung tautly; upon its surface hung a fragrant plum blossom. A poem also recited thus spake: "For twenty years I have sought truth amidst strife, where blooming durian flowers illuminate the palace walls. In spring’s embrace, they vie for beauty, tigers clashing in dreams of return." MT 0.60
ZS-Exppt And as he turned his gaze back upon the scene once more, there was displayed an archer’s bow strung tautly; upon its surface hung a fragrant plum blossom. A poem also recited thus: "For twenty years I have sought truth amidst strife, where blooming durian flowers illuminate the palace walls. In spring’s embrace, they vie for beauty, tigers clashing in dreams of return." MT 0.61
@@
Table 7: Example paragraphs from “Dream of the Red Chamber” (ZH $\rightarrow$ EN) translated with H1personalization. Setup outputs are from Gemma 2 2B; only MS and SAE Cont.pt successfully flip the [Uncaptioned image] classifier’ prediction.
@@
ZS-Expht Meanwhile, the stable door was shut, leaving Pinocchio alone. His tummy rumbled because it had been hours since his last meal, and he started yawning widely from his hunger. With each yawn, his mouth opened like an oven. MT 0.78
ZS-Exppt Meanwhile, the stable door was shut, leaving Pinocchio all alone. His tummy rumbled because it had been hours since his last meal, and he started yawning widely from sheer hunger. With each yawn, his mouth gaped open like an oven. H2 0.80
@@
Table 8: Example paragraphs from “Pinocchio” (IT $\rightarrow$ EN) translated with H2personalization. Setup outputs are from Gemma 2 9B; ZS-Exppt, MS, SAE Cont.ht and SAE Cont.pt (both $\alpha=5$) can flip the [Uncaptioned image] classifier prediction.

### v2

@@ Appendix A Experimental Reproducibility
A.1 Base prompt
We use the same prompt template across all methods: ZS (which corresponds to the original model translation), ZS-Exp. (detailed in Section˜A.5), MS, ActAdd, ReFT, and SAE Cont. This prompt, shown in Listing LABEL:lst:explain-final, instructs the model to translate the source sentence while explicitly preventing it from adding any explanations about the translation process. Since all test models are Instruction Tuned, we utilize their native chat templates to preprocess the input accordingly. For multi-shot examples, the user and assistant turns are repeated for each example, always using the same prompt structure.
A.2 Classifier training
@@
Training data only includes generations from models and the translator without any source text. It is also perfectly balanced, as each paragraph provides one instance for all three labels: H1, H2, and MT. The total size of the training set varies depending on the number of paragraphs in the chosen novel. On average, we obtain approximately 830 instances, resulting in a total of around 2,490 labeled examples for training (see 1). Validation and test sets are strictly held out and never seen during training. Additionally, they do not include the small 20-example subsets used in the MS, ZS-Exp. and SAE Cont. setups.
A.3 ReFT training
ReFT training was conducted using the PyReFT toolkit from the original authors121212stanfordnlp/pyreft. We applied the intervention at the same hook point used by other steering methods - specifically, the layer output corresponding to the residual stream at the selected layer. The training configuration includes a low rank dimension of 4, lora alpha set to 32, and a lora dropout of 0.05. ReFT was trained on the same 20 prompts used in the MS setup, for a total of 100 epochs.
A.4 SAE Cont.
@@ A.3 ReFT training
A.5 ZS-Exp.
For ZS-Exp., we used GPT-4o (June 2025) to generate explanations detailing the stylistic differences between a base translation and a target human translation. The prompt template used for this task is shown in Listing LABEL:lst:explain-template, using the same 20 examples as in the MS and SAE Cont. setups. Example outputs for different novels are shown in Listing LABEL:lst:explain-example. Finally, these generated guidelines are used to prompt the evaluated models, following the template shown in Listing LABEL:lst:explain-final.
@@
Base translation: <MT text here>
@@
Appendix B Results across languages
B.1 Classifiers
@@
B.2 Prompting and steering results
@@
Appendix C Examples from Dataset and Conditioning Outputs
We present in Tables 8 and 9 a selection of examples from two different languages, showcasing outputs from each of the tested setups. For each example, we also report the corresponding classification label predicted by the [Uncaptioned image] classifier and the associated [Uncaptioned image]  Comet score. Additionly table 10 shows some examples of models generating output aligned with the Human translator according to the [Uncaptioned image] classifier but with a low [Uncaptioned image]  Comet score corresponding to an almost unreadable output due to extreme $\alpha$ values.
Appendix D Additional Experiment on Contrastive Setups to Isolate Style Features
To isolate features relevant to a specific translation style, we adopt a contrastive approach between different parallel translation examples. Specifically we test two setups:
  • $M\leftrightarrow H$: contrasting the target style we aim to achieve ($H^{+}$) with a "clean" style derived from the model’s original machine translation ($M$);
  • $H_{\alpha}\leftrightarrow H_{\beta}$ contrasting the desired style with the style of another human translator.
The $M\leftrightarrow H$ setup is designed to extract features that distinguish the target human-like style from a standard, model-generated translation. The $H_{\alpha}\leftrightarrow H_{\beta}$ setup, on the other hand, aims to isolate the features that differentiate one translator’s style from another. While the latter more closely corresponds to the notion of translation style (as a property unique to each translator), we chose to experiment with both since prior works have shown that contrasting the desired behavior with the model’s baseline behavior tends to yield the most effective results when steering LLMs.
Which contrastive setup is better?
We present in Table 7 the results for both contrastive setups. For both the ZS-Exp. and SAE Cont. approaches, the results are generally comparable, making it difficult to identify a clear best-performing method. However, larger models occasionally achieve higher performance, which we hypothesize may stem from their enhanced ability to disentangle personalization-relevant features even without explicit guidance.
Appendix E Additional Results when Scaling Demonstrations Count
Figure˜5 presents the full set of experiments, including two languages (Russian and French) and two models (Gemma 2 2B and 9B), complementing the results of Figure˜4. The two languages were selected as they presented the largest collection of translated paragraphs available (see Table˜2), allowing us to scale the ICL set without substantially reducing the size of the training data or impairing classifier performance. Notably, as also shown in Figure˜4 the personalization accuracy P of our contrastive approach improves across all models and language tested, even surpassing the multi-shot (MS) baseline. In all cases, the MS baseline reaches the LLM context size limit (around ICL $\geq 50$), after which the performance degrades significantly. Similarly, the tok/s efficiency metric highlights the advantage of our contrastive approach: It brings only a minimal, roughly linear, decrease in inference-time speed (compared to ZS, i.e. ICL =0=0), whereas the MS setup exhibits a progressively larger computational impact as the number of in-context examples increases.
Figure 5: Complete results when comparing the MS approach to our SAE Cont.M↔H for the Gemma models (2B and 9B) on the largest novels, evaluated at the paragraph level, in Russian (RU) and French (FR).
@@ Appendix E Additional Results when Scaling Demonstrations Count
Method H P $_{\textsc{flip}}$ [Uncaptioned image] tok/s
Gemma 2 2B
ZS-ExpM↔H 0.30 0.22 0.16 0.68 41.3
[…]
SAE Cont.M↔H 0.39 0.27 0.19 0.70 41.1
[…]
Gemma 2 9B
ZS-ExpM↔H 0.41 0.22 0.18 0.72 24.6
[…]
SAE Cont.M↔H 0.46 0.33 0.29 0.72 25.1
[…]
LLaMA 3.1 8B
ZS-ExpM↔H 0.56 0.23 0.21 0.69 25.5
[…]
SAE Cont.M↔H 0.59 0.31 0.27 0.72 23.1
[…]
Table 7: Averaged metric scores across all tested languages for the $M\leftrightarrow H$ and $H_{\alpha}\leftrightarrow H_{\beta}$ variants of ZS-Exp. and SAE Cont. methods. H: human style accuracy, i.e. p(p(H1)+p()+p(H2)). P: personalization accuracy $p(H_{x})$ for the target style. $_{\textsc{flip}}$: Proportion of segments for which steering has a causal impact on personalization. $\alpha=5$ is used for SAE Cont. results.
@@
ZS-ExpM↔H And as he turned his gaze back upon the scene once more, there was displayed an archer’s bow strung tautly; upon its surface hung a fragrant plum blossom. A poem also recited thus spake: "For twenty years I have sought truth amidst strife, where blooming durian flowers illuminate the palace walls. In spring’s embrace, they vie for beauty, tigers clashing in dreams of return." MT 0.60
[…]
@@
[…]
@@
ZS-ExpM↔H Meanwhile, the stable door was shut, leaving Pinocchio alone. His tummy rumbled because it had been hours since his last meal, and he started yawning widely from his hunger. With each yawn, his mouth opened like an oven. MT 0.78
[…]
@@
[…]